# An extension of Viennot’s shadow to rook placements via orbit harmonics

### v2

@@ Abstract.
For fixed positive integers $n,m,r$ with $r\leq\min(n,m)$, let ${\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ be the affine space of $n\times m$ complex matrices with coordinate ring ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$. We define a homogeneous ideal $I_{n,m,r}$, where the graded quotient ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ carries an action of the group ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ by independent permutations of rows and columns. This quotient ring is obtained by applying the orbit harmonics method to matrix loci corresponding to all rook placements of size at least $r$. We show that the quotient ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ admits a standard monomial basis, which is determined by extending rook placements to an element in ${\mathfrak{S}}_{n+m-r}$ and applying Viennot’s shadow line avatar of the Schensted correspondence. We also determine the graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module structure of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$.
@@ 1. Introduction
Let ${\mathbf{x}}_{N}=(x_{1},x_{2},\dots,x_{N})$ be a collection of $N$ variables, and let ${\mathbb{C}}[{\mathbf{x}}_{N}]:={\mathbb{C}}[x_{1},x_{2},\dots,x_{N}]$ be the polynomial ring over these variables. For a finite locus of points ${\mathcal{Z}}\subseteq{\mathbb{C}}^{N}$, its vanishing ideal is
@@
The coordinate ring ${\mathbb{C}}[{\mathcal{Z}}]$ can then be identified with ${\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathbf{I}}({\mathcal{Z}})$. The method of orbit harmonics replaces the ideal ${\mathbf{I}}({\mathcal{Z}})$ with its associated graded ideal ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}})$, extending the identification of (ungraded) vector spaces to
@@
Geometrically, the method of orbit harmonics linearly deforms the locus $\mathcal{Z}$} to a scheme of multiplicity $|{\mathcal{Z}}|$ supported at the origin. In the picture below, this deformation is drawn schematically for $\mathcal{Z}$}, which is a locus of six points stable under the action of ${\mathfrak{S}}_{3}$ generated by reflections across the three lines.
@@
In this paper, we consider a matrix locus formed by rook placements. For positive integers $n$ and $m$, a rook placement $\mathcal{R}$} on the $n\times m$ board is a subset of $n]\times[m]$ such that $(i,j)\in{\mathcal{R}}$ and $(i^{\prime},j^{\prime})\in{\mathcal{R}}$ implies $i\neq i^{\prime}$ and $j\neq j^{\prime}$. We call each element $(i,j)\in{\mathcal{R}}$ a rook of $\mathcal{R}$}. Define the size $|{\mathcal{R}}|$ of $\mathcal{R}$} by
@@
The locus ${\mathcal{UZ}}_{n,m,r}$ is closed under the action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ by left and right multiplication, and the quotient $R({\mathcal{UZ}}_{n,m,r})$ is a graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module. Our results on this quotient include:
@@ 2.2. Orbit Harmonics
Let ${\mathcal{Z}}\subset{\mathbb{C}}^{N}$ be a finite locus of points in the affine $N$-space. The vanishing ideal ${\mathbf{I}}({\mathcal{Z}})\subseteq{\mathbb{C}}[{\mathbf{x}}_{N}]$ is the ideal
@@
Given a set of generators $I=\langle f_{1},f_{2},\dots,f_{r}\rangle$, we have the containment of ideals $\langle\tau(f_{1}),\tau(f_{2}),\dots,\tau(f_{r})\rangle\subseteq{\mathrm{gr}}\,I$. This containment is strict in general. On the other hand, if $\{f_{1},f_{2},\dots,f_{r}\}$ forms a Gröbner basis of $I$ with respect to some graded monomial order, we have the equality $\langle\tau(f_{1}),\tau(f_{2}),\dots,\tau(f_{r})\rangle={\mathrm{gr}}\,I$.
@@ 2.4. Representation Theory
We introduce the truncation operator $\{-\}_{P}$ where $P$ is a condition. Given $F=\sum_{\lambda}c_{\lambda}s_{\lambda}\in\Lambda$ with $c_{\lambda}\in{\mathbb{C}}(q)$, we define $\{F\}_{P}$ by
@@
Irreducible representations of ${\mathfrak{S}}_{n}$ are indexed by partitions $\lambda\vdash n$. We write $V^{\lambda}$ for the irreducible ${\mathfrak{S}}_{n}$-module indexed by $\lambda$. Given an ${\mathfrak{S}}_{n}$-module $V$, there exists a unique decomposition $V\cong\bigoplus_{\lambda\vdash n}c_{\lambda}V^{\lambda}$ with $c_{\lambda}\geq 0$. The Frobenius image of $V$ is defined to be
(2.8) ${\mathrm{Frob}}(V)=\sum_{\lambda\vdash n}c_{\lambda}s_{\lambda},$
@@
(2.9) ${\mathrm{grFrob}}(V,q)=\sum_{d\geq 0}{\mathrm{Frob}}(V_{d})\cdot q^{d}.$
@@
(2.10) ${\mathrm{Frob}}(V)=\sum_{\begin{subarray}{c}\lambda\vdash n\\ \mu\vdash m\end{subarray}}c_{\lambda,\mu}s_{\lambda}\otimes s_{\mu},$
and define the graded Frobenius images for graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ analogously.
@@
For any irreducible ${\mathfrak{S}}_{n}$-module $V^{\lambda}$, we have
@@
(2.13) $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}=\bigoplus_{\begin{subarray}{c}\mu\vdash j,\\ \mu_{i}\leq\lambda_{i}\text{ for all }i\end{subarray}}V^{\mu},$
which implies that $V^{(j)}$ has positive multiplicity in $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}$ if and only if $\lambda_{1}\geq j$, completing the proof. ∎
@@ 3. The ideal $I_{n,m,r}$ and its structure
  • Every entry in the matrix is 0 or 11
  • There is at most one nonzero entry in each row or column
@@
Our object of study is the upper rook placements, defined as
@@
the union of all rook placements of size at least $r$. To study the orbit harmonics quotient $R({\mathcal{Z}}_{n,m,r})$, we define the following ideal, which we later prove to be ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$.
@@
  • any product $\prod_{k=1}^{n-r+1}\big(\sum_{j=1}^{m}x_{i_{k},j}\big)$ of nr+1n-r+1 distinct row sums for $1\leq i_{1}<\cdots<i_{n-r+1}\leq n$, and
  • any product $\prod_{k=1}^{m-r+1}\big(\sum_{i=1}^{n}x_{i,j_{k}}\big)$ of mr+1m-r+1 distinct column sums for $1\leq j_{1}<\cdots<j_{m-r+1}\leq m$.
@@ 3.1. $r=0$ case
For convenience, write ${\mathcal{Z}}_{n,m}={\mathcal{UZ}}_{n,m,0}$ for the set of all rook placements on the $n\times m$ board. Let $I_{n,m}=I_{n,m,0}$ be the ideal generated by
  • any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $1\leq i\leq n$ and $1\leq j,j^{\prime}\leq m$ of variables in the same row, and
@@
We start by giving a spanning set of the vector space ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$:
@@
We now compare dimensions. Lemma 3.3 indicates that
@@
It suffices to study the module structure of ${\mathbb{C}}[{\mathcal{Z}}_{n,m,d}]$ for $0\leq d\leq\min\{m,n\}$. Consider ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d}$ (resp. ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d}$) as a subgroup of ${\mathfrak{S}}_{n}$ (resp. ${\mathfrak{S}}_{m}$) by letting ${\mathfrak{S}}_{d}$ and ${\mathfrak{S}}_{n-d}$ (resp. ${\mathfrak{S}}_{m-d}$) separately act on $\{1,\dots,d\}$ and $\{d+1,\dots,n\}$ (resp. $\{d+1,\cdots,m\}$). Thus, we embed $({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})$ into ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ as a subgroup. Note that $({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})$ acts on ${\mathbb{C}}[{\mathfrak{S}}_{d}]$ by
@@
$\cong\mathrm{Ind}_{({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})}^{{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}}\bigg(\bigoplus_{\mu\vdash d}(V^{\mu}\otimes V^{n-d})\otimes(V^{\mu}\otimes V^{m-d})\bigg)$
$\cong\bigoplus_{\mu\vdash d}\mathrm{Ind}_{({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})}^{{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}}\big((V^{\mu}\otimes V^{n-d})\otimes(V^{\mu}\otimes V^{m-d})\big)$
$\cong\bigoplus_{\mu\vdash d}\Big(\mathrm{Ind}_{{\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d}}^{{\mathfrak{S}}_{n}}(V^{\mu}\otimes V^{n-d})\Big)\otimes\Big(\mathrm{Ind}_{{\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d}}^{{\mathfrak{S}}_{m}}(V^{\mu}\otimes V^{m-d})\Big)$
@@ 3.2. Structure of $I_{n,m,r}$
In order to analyze the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ in Section 4 and the graded module structure of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ in Section 5, we start by studying strategic elements of the ideal $I_{n,m,r}$. On a first reading, it may be better to skip all the technical proofs in this section and accept Corollary 3.9 and Corollary 3.10 on faith.
@@
  • all products $\prod_{k=1}^{m-r+1}\big(\sum_{i=1}^{n}x_{i,j_{k}}\big)$ of mr+1m-r+1 distinct column sums for $1\leq j_{1}<\cdots<j_{m-r+1}\leq m$.
@@
If $m-r<t\leq m$ and $b_{1},\cdots,b_{t}\in[m]$ are distinct integers, then
$\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t}\in[n]\\ \text{distinct}\end{subarray}}\prod_{j=1}^{t}x_{i_{j},b_{j}}\in J_{n,m,r}$
@@
$\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t}\in[n]\\ \text{distinct}\end{subarray}}\prod_{j=1}^{t}x_{i_{j},b_{j}}\equiv\sum_{i_{1},\cdots,i_{t}\in[n]}\prod_{j=1}^{t}x_{i_{j},b_{j}}=\prod_{j=1}^{t}\sum_{i=1}^{n}x_{i,b_{j}}\mod{J_{n,m,r}}.$
@@
Write ${\mathfrak{m}}({\mathcal{R}})=\prod_{k=1}^{d}x_{a_{k},b_{k}}$ where $b_{1},\cdots,b_{d}\in[m]$ are distinct and $1\leq a_{1}<\cdots<a_{d}\leq n$. Since $p+d>n+m-r\geq n$, it follows that $p]\cap\{a_{1},\cdots,a_{d}\}\neq\varnothing$. Therefore, we can choose the largest $t\in[d]$ such that $a_{t}\in[p]$.
@@
In fact, the maximality of $t$ implies that $p]\cap\{a_{t+1},\cdots,a_{d}\}=\varnothing$, so we have that $p]\sqcup\{a_{t+1},\cdots,a_{d}\}\subseteq[n]$ and hence $p+d-t\leq n$. Consequently, $t\geq p+d-n>m-r$, which completes the proof of the claim.
@@
$(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})=(\eta_{p}\otimes 1)\cdot\prod_{k=1}^{d}x_{a_{k},b_{k}}=(p-t)!\cdot\Bigg(\prod_{\begin{subarray}{c}i_{1},\cdots,i_{t}\in[p]\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}$
$\equiv$ $(p-t)!\cdot\Bigg(-\sum_{\begin{subarray}{c}\text{distinct $i_{1},\cdots,i_{t}\in[n]$,}\\ \{i_{1},\cdots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}$
$\equiv$ $(p-t)!\cdot\Bigg(-\sum_{\begin{subarray}{c}\text{distinct $i_{1},\cdots,i_{t}\in[n]\setminus\{a_{t+1},\cdots,a_{d}\}$,}\\ \{i_{1},\cdots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\mod{J_{n,m,r}}$

### v3

@@ Abstract.
For fixed positive integers $n,m,r$, let ${\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ be the affine space of $n\times m$ complex matrices with coordinate ring ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$. We define a homogeneous ideal $I_{n,m,r}$, where the graded quotient ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ is obtained from the orbit harmonics deformation of the matrix loci corresponding to all rook placements of size at least $r$. By extending rook placements to elements in ${\mathfrak{S}}_{n+m-r}$ and applying Viennot’s shadow line avatar of the Schensted correspondence, we compute the standard monomial basis of the quotient ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ with respect to diagonal monomial orders. We also determine the graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module structure of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$.
@@ 1. Introduction
Let ${\mathbf{x}}_{N}=(x_{1},x_{2},\dots,x_{N})$ be a collection of $N$ variables and let ${\mathbb{C}}[{\mathbf{x}}_{N}]:={\mathbb{C}}[x_{1},x_{2},\dots,x_{N}]$ be the polynomial ring over these variables. For a finite locus of points ${\mathcal{Z}}\subseteq{\mathbb{C}}^{N}$, its vanishing ideal is
@@
The coordinate ring ${\mathbb{C}}[{\mathcal{Z}}]$ can then be identified with ${\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathbf{I}}({\mathcal{Z}})$. The method of orbit harmonics replaces the ideal ${\mathbf{I}}({\mathcal{Z}})$ with its associated graded ideal ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}})$, extending this identification of (ungraded) vector spaces to
@@
Geometrically, the method of orbit harmonics linearly deforms the locus $\mathcal{Z}$} to a scheme of multiplicity $|{\mathcal{Z}}|$ supported at the origin. In the picture below, this deformation is drawn schematically for $\mathcal{Z}$}, where $\mathcal{Z}$} is a locus of six points stable under the action of ${\mathfrak{S}}_{3}$ generated by reflections across the three lines.
@@
In this paper, we consider a matrix locus formed by rook placements. For any positive integer $n$, let $[n]$ be the set of integers $\{1,2,\dots,n\}$. For positive integers $n$ and $m$, a rook placement $\mathcal{R}$} on the $n\times m$ board is a subset of $n]\times[m]$ such that $(i,j)\in{\mathcal{R}}$ and $(i^{\prime},j^{\prime})\in{\mathcal{R}}$ implies $i\neq i^{\prime}$ and $j\neq j^{\prime}$. We call each element $(i,j)\in{\mathcal{R}}$ a rook of $\mathcal{R}$}. Define the size $|{\mathcal{R}}|$ of $\mathcal{R}$} by
@@
The locus ${\mathcal{UZ}}_{n,m,r}$ is closed under the action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ by row and column permutation, and the quotient $R({\mathcal{UZ}}_{n,m,r})$ is a graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module. Our results on this quotient include:
@@ 2.2. Orbit Harmonics
Let ${\mathcal{Z}}\subset{\mathbb{C}}^{N}$ be a finite locus of points in affine $N$-space. The vanishing ideal ${\mathbf{I}}({\mathcal{Z}})\subseteq{\mathbb{C}}[{\mathbf{x}}_{N}]$ is the ideal
@@
Given a set of generators $I=\langle f_{1},f_{2},\dots,f_{r}\rangle$, we have the containment of ideals $\langle\tau(f_{1}),\tau(f_{2}),\dots,\tau(f_{r})\rangle\subseteq{\mathrm{gr}}\,I$. This containment is strict in general. On the other hand, if << is a monomial order such that $f\leq g$ if and only if $\deg(f)\leq deg(g)$ and $\{f_{1},f_{2},\dots,f_{r}\}$ forms a Gröbner basis of $I$ with respect to <<, we have the equality $\langle\tau(f_{1}),\tau(f_{2}),\dots,\tau(f_{r})\rangle={\mathrm{gr}}\,I$.
@@ 2.4. Representation Theory
We introduce the truncation operator $\{-\}_{P}$ where $P$ is a condition that partitions may or may not satisfy. Given $F=\sum_{\lambda}c_{\lambda}s_{\lambda}\in\Lambda$ with $c_{\lambda}\in{\mathbb{C}}(q)$, we define $\{F\}_{P}$ by
@@
Irreducible representations of ${\mathfrak{S}}_{n}$ are indexed by partitions $\lambda\vdash n$. We write $V^{\lambda}$ for the irreducible ${\mathfrak{S}}_{n}$-module indexed by $\lambda$. Given an ${\mathfrak{S}}_{n}$-module $V$, there exists a unique decomposition $V\cong\bigoplus_{\lambda\vdash n}c_{\lambda}V^{\lambda}$ with $c_{\lambda}\geq 0$. The Frobenius image of $V$ is defined to be the symmetric function
(2.8) ${\mathrm{Frob}}(V):=\sum_{\lambda\vdash n}c_{\lambda}s_{\lambda},$
@@
(2.9) ${\mathrm{grFrob}}(V,q):=\sum_{d\geq 0}{\mathrm{Frob}}(V_{d})\cdot q^{d}.$
@@
(2.10) ${\mathrm{Frob}}(V):=\sum_{\begin{subarray}{c}\lambda\vdash n\\ \mu\vdash m\end{subarray}}c_{\lambda,\mu}s_{\lambda}\otimes s_{\mu},$
and define the graded Frobenius images for graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-modules analogously.
@@
For any ${\mathfrak{S}}_{n}$-module $V^{\lambda}$, we have
@@
(2.13) $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{n-1}}\,V^{\lambda}=\bigoplus_{\begin{subarray}{c}\mu\vdash n-1,\\ \mu_{i}\leq\lambda_{i}\text{ for all }i\end{subarray}}V^{\mu},$
iterating which implies that $V^{(j)}$ has positive multiplicity in $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}$ if and only if $\lambda_{1}\geq j$, completing the proof. ∎
@@ 3. The ideal $I_{n,m,r}$ and its structure
  • every entry in the matrix is 0 or $1$,
  • there is at most one nonzero entry in each row or column,
@@
The locus ${\mathcal{Z}}_{n,m,r}$ itself is hard to analyze using orbit harmonics. However, it turns out that the locus of upper rook placements, defined as
@@
the union of all rook placements of size at least $r$, behaves nicely with the method of orbit harmonics. To study the orbit harmonics quotient $R({\mathcal{Z}}_{n,m,r})$, we define the following ideal, which we later prove to be ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$.
@@
  • any product $\prod_{k=1}^{n-r+1}\big(\sum_{j=1}^{m}x_{i_{k},j}\big)$ of nr+1n-r+1 distinct row sums for $1\leq i_{1}<\dots<i_{n-r+1}\leq n$, and
  • any product $\prod_{k=1}^{m-r+1}\big(\sum_{i=1}^{n}x_{i,j_{k}}\big)$ of mr+1m-r+1 distinct column sums for $1\leq j_{1}<\dots<j_{m-r+1}\leq m$.
@@ 3.1. $r=0$ case
When $r=0$, write ${\mathcal{Z}}_{n,m}:={\mathcal{UZ}}_{n,m,0}$ for the locus of all rook placements on the $n\times m$ board. The locus ${\mathcal{Z}}_{n,m}$ is a finite shifted locus, which was defined and studied by Reiner and Rhoades [12, Sec. 3.4]. Lemma 3.3 and Proposition  3.4 below can be derived directly from the results by Reiner and Rhoades, but we still include their proofs as they will promote understanding of later sections.
Write $I_{n,m}:=I_{n,m,0}$ for the ideal generated by
  • any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $1\leq i\leq n$ and $1\leq j,j^{\prime}\leq m$ of variables in the same row,
@@
Note that the products of row sums and column sums are no longer present as it’s impossible to have $n+1$ distinct rows or $m+1$ distinct columns. We start by giving a spanning set of the vector space ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$:
@@
We now compare dimensions. Lemma 3.3 implies that
@@
It suffices to study the module structure of ${\mathbb{C}}[{\mathcal{Z}}_{n,m,d}]$ for $0\leq d\leq\min\{m,n\}$. Consider ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d}$ (resp. ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d}$) as a subgroup of ${\mathfrak{S}}_{n}$ (resp. ${\mathfrak{S}}_{m}$) by letting ${\mathfrak{S}}_{d}$ and ${\mathfrak{S}}_{n-d}$ (resp. ${\mathfrak{S}}_{m-d}$) separately act on $\{1,\dots,d\}$ and $\{d+1,\dots,n\}$ (resp. $\{d+1,\dots,m\}$). Thus, we embed $({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})$ into ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ as a subgroup. Note that $({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})$ acts on ${\mathbb{C}}[{\mathfrak{S}}_{d}]$ by
@@
$\cong\mathrm{Ind}_{({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})}^{{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}}\bigg(\bigoplus_{\mu\vdash d}(V^{\mu}\otimes V^{(n-d)})\otimes(V^{\mu}\otimes V^{(m-d)})\bigg)$
$\cong\bigoplus_{\mu\vdash d}\mathrm{Ind}_{({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})}^{{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}}\big((V^{\mu}\otimes V^{(n-d)})\otimes(V^{\mu}\otimes V^{(m-d)})\big)$
$\cong\bigoplus_{\mu\vdash d}\Big(\mathrm{Ind}_{{\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d}}^{{\mathfrak{S}}_{n}}(V^{\mu}\otimes V^{(n-d)})\Big)\otimes\Big(\mathrm{Ind}_{{\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d}}^{{\mathfrak{S}}_{m}}(V^{\mu}\otimes V^{(m-d)})\Big)$
@@ 3.2. Structure of $I_{n,m,r}$
In order to analyze the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ in Section 4 and the graded module structure of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ in Section 5, we start by studying strategic elements of the ideal $I_{n,m,r}$. On a first reading, it may be better to skip all the technical proofs in this section and accept Corollary 3.9 on faith.
@@
  • all products $\prod_{k=1}^{m-r+1}\big(\sum_{i=1}^{n}x_{i,j_{k}}\big)$ of mr+1m-r+1 distinct column sums for $1\leq j_{1}<\dots<j_{m-r+1}\leq m$.
@@
If $m-r<t\leq m$ and $b_{1},\dots,b_{t}\in[m]$ are distinct integers, then
$\sum_{\begin{subarray}{c}i_{1},\dots,i_{t}\in[n]\\ \text{distinct}\end{subarray}}\prod_{j=1}^{t}x_{i_{j},b_{j}}\in J_{n,m,r}$
@@
$\sum_{\begin{subarray}{c}i_{1},\dots,i_{t}\in[n]\\ \text{distinct}\end{subarray}}\prod_{j=1}^{t}x_{i_{j},b_{j}}\equiv\sum_{i_{1},\dots,i_{t}\in[n]}\prod_{j=1}^{t}x_{i_{j},b_{j}}=\prod_{j=1}^{t}\sum_{i=1}^{n}x_{i,b_{j}}\mod{J_{n,m,r}}.$
@@
Write ${\mathfrak{m}}({\mathcal{R}})=\prod_{k=1}^{d}x_{a_{k},b_{k}}$ where $b_{1},\dots,b_{d}\in[m]$ are distinct and $1\leq a_{1}<\dots<a_{d}\leq n$. Since $p+d>n+m-r\geq n$, it follows that $p]\cap\{a_{1},\dots,a_{d}\}\neq\varnothing$. Therefore, we can choose the largest $t\in[d]$ such that $a_{t}\in[p]$.
@@
In fact, the maximality of $t$ implies that $p]\cap\{a_{t+1},\dots,a_{d}\}=\varnothing$, so we have that $p]\sqcup\{a_{t+1},\dots,a_{d}\}\subseteq[n]$ and hence $p+d-t\leq n$. Consequently, $t\geq p+d-n>m-r$, which completes the proof of the claim.
@@
$(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})=(\eta_{p}\otimes 1)\cdot\prod_{k=1}^{d}x_{a_{k},b_{k}}=(p-t)!\cdot\Bigg(\sum_{\begin{subarray}{c}i_{1},\dots,i_{t}\in[p]\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\dots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}$
$\equiv$ $(p-t)!\cdot\Bigg(-\sum_{\begin{subarray}{c}\text{distinct $i_{1},\dots,i_{t}\in[n]$,}\\ \{i_{1},\dots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\dots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}$
$\equiv$ $(p-t)!\cdot\Bigg(-\sum_{\begin{subarray}{c}\text{distinct $i_{1},\dots,i_{t}\in[n]\setminus\{a_{t+1},\dots,a_{d}\}$,}\\ \{i_{1},\dots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\dots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\mod{J_{n,m,r}}$